OpenStudy (mtalhahassan2):

For each of the following functions, calculate the average rate of change on the interval XE [2,5]. a) f(x)= 3x+1

OpenStudy (matt101):

Since this is a straight line, the slope of that line (its rate of change) is constant no matter what interval you look at. The equation is given to you in y=mx+b form, where m is the slope. So, using that information...what is the average rate of change of this function?

OpenStudy (mtalhahassan2):

but how can we solve this question!!

OpenStudy (mtalhahassan2):

Plz help me

OpenStudy (mtalhahassan2):

@razor99

OpenStudy (k8lyn911):

y=mx+b y=3x+1 m = 3

OpenStudy (mtalhahassan2):

so basically we have to find the slope

OpenStudy (k8lyn911):

Yes. Rate of change = slope.

OpenStudy (mtalhahassan2):

but how can we find for this one g(x)= 1/x

OpenStudy (mtalhahassan2):

@k8lyn911

OpenStudy (anonymous):

yes k8 is right

OpenStudy (mtalhahassan2):

@ Kevindragonlord

OpenStudy (mtalhahassan2):

@Kevindragonlord

OpenStudy (matt101):

Well g(x)=1/x is a curved line, so the slope changes depending on where on the line you're looking. If you want to find the slope at a particular point on the line you need calculus, but if all you need is the average slope between two points on that line, just calculate the slope using the equation for slope: \[m={y_2-y_1 \over x_2-x_1}\]